OpenStudy (anonymous):

A particle projected from ground takes time 't1' to reach 15/16 H height and time 't2' to cover rest of distance to the ground. If 'H' is the maximum height attaine. please find the ratio t1: t2

OpenStudy (anonymous):

@Abhisar @robtobey

OpenStudy (anonymous):

@dan815 @Destinymasha @mathmale @zzr0ck3r

OpenStudy (anonymous):

@uri @paki @Whitemonsterbunny17

OpenStudy (anonymous):

projected from the ground then \[X = \frac{ 1 }{ 2 } a t ^{2}\] \[t=\sqrt{\frac{ 2 X }{ a }}.\] x for t1 = 15/16 H x for t2 = 17/16 H Thus divide them and symbols will be cancelled.

OpenStudy (anonymous):